Calculate the pressure exerted by 0.0153 moles of CO2 gas in a container of 1.467 L at 20.0°C.  

1
Expert's answer
2021-03-01T02:35:14-0500

PV=nRT

P×1,467=0.0153×8.31×293

P×1,467=37.25

P=25,39